OpenStudy (anonymous):

2log50=3 log 25+log(x-2) helppppp please

OpenStudy (anonymous):

log(a) + log(b) = log(ab) log(c) - log(d) = log(c/d)

OpenStudy (anonymous):

and also alog(b) = log(b^a)

OpenStudy (anonymous):

2log50=3 log 25+log(x-2) 2log50-3 log 25=log(x-2) now use log(c) - log(d) = log(c/d)

OpenStudy (anonymous):

your job is to write \[\log(x-2)=\log(whatever)\] so you can set \[x-2=whatever\] and solve

OpenStudy (anonymous):

before that use alog(b) = log(b^a)

OpenStudy (anonymous):

The key to this problem is simplifying the left hand side after you isolate the term with the x on the right.\[2\log50=3 \log 25+\log(x-2) \implies 2\log50-3 \log 25=\log(x-2)\]\[\log50^2- \log 25^3=\log(x-2) \implies \log(\frac{50^2}{ 25^3})=\log(\frac{4}{25})=\log(x-2)\]\[\implies \frac{4}{25}=x-2 \implies \frac{54}{25}=x\]

OpenStudy (anonymous):

If you don't simplify, the numbers in the fraction get discouragingly large.